OpenStudy (anonymous):

general form of the equation of Center (0,0) radius 3

OpenStudy (lalaly):

x^2 +y^2 = 9

OpenStudy (anonymous):

\[x^2 + y^2 = r^2 \]\[x^2 + y^2 = 9\]

OpenStudy (anonymous):

what if the center is changed?

OpenStudy (anonymous):

\[(x -h)^2 +(y-k)^2 = r^2\] \[(h,k) \text{ is the center}\]

OpenStudy (anonymous):

but thats standard form

OpenStudy (anonymous):

oh I see

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

do i have to put it in standard form then change it? or is there another way

OpenStudy (anonymous):

\[x^2 + y^2 +2gy +2hx +c=0\] \[(-h,-g)\text{ is the center }\]